OpenStudy (anonymous):

Suppose that triangle ABC is similar to triangle DEF with AB = 20, BC = 8, AC = 19, and EF = 12. Find DE and DF.

Directrix (directrix):

Lengths of corresponding sides of similar triangles are in proportion. The similarity statement triangle ABC is similar to triangle DEF gives that AB/DE = BC/EF = AC/DF. So, 20/DE = 8/12 = 19/DF The scale factor is 8/12 or 2/3. 20/DE = 8/12 8 (DE) = 12*20 DE = 12*(20/8) DE = 30 --------- Your turn now: Find DF by solving: 8/12 = 19/DF Post answer here for free check.

OpenStudy (anonymous):

27=df

Directrix (directrix):

@1dpp --> Check my work and yours, too. 19/DF = 8/12 8 DF = 19*12 DF = (19*12) / 8 DF = 28.5

OpenStudy (anonymous):

I saw what i did. i made a silly mistake and multiplied 18*12 smh.. Thanks